OpenStudy (anonymous):

A 71 kg man weighs more at the north pole than at the equator. By how much?

OpenStudy (anonymous):

update: tried 3.692 and it was incorrect

OpenStudy (anonymous):

You would weigh about 0.3% less at the equator than at the poles.

OpenStudy (michele_laino):

weight of an object depends on the gravity. So we need to know what is gravity at North Pole and what is the gravity at Equator

OpenStudy (anonymous):

On wikipedia it says gravity is 9.780 m/s^2 at the equator and 9.832 m/s^2 at the north pole

OpenStudy (mrnood):

There is also a very minor effect of rotation - the body at the pole is not travelling in the circle around the equator. Since the period is 24hours I would suspect this would be an almost insignificant effect....

OpenStudy (michele_laino):

then the weight is 70*9.832=...?Newtons at North pole and 70*9.780=... Newtons at Equator

OpenStudy (anonymous):

71 you mean?

OpenStudy (michele_laino):

yes! sorry!

OpenStudy (anonymous):

I did both and subtracted the difference and got 3.962 N

OpenStudy (mrnood):

hmmm - my rough calculation makes the rotation effect approx. 2N so maybe significant

OpenStudy (michele_laino):

no, you have to compute the ratio, between the weight at north pole over the weight at Equator, like this: \[\Large \frac{{71 \times 9.832}}{{71 \times 9.780}} = \frac{{9.832}}{{9.870}} = ...?\]

OpenStudy (michele_laino):

oops: \[\Large \frac{{71 \times 9.832}}{{71 \times 9.780}} = \frac{{9.832}}{{9.780}} = ...?\]

OpenStudy (mrnood):

@Michele_Laino It asks for the difference in weight - not the ratio subtract the 2 values....

OpenStudy (michele_laino):

yes! you are right! @MrNood

OpenStudy (anonymous):

I got 3.962 and it's incorrect

OpenStudy (anonymous):

Does someone mind checking my math?

OpenStudy (mrnood):

your maths is correct if the figures are correct

OpenStudy (mrnood):

How do you know your answer is incorrect? what value do you have for an answer?

OpenStudy (anonymous):

its an online program that gives immediate feedback if I submit a wrong answer. It's in newtons

OpenStudy (mrnood):

so you have no ide if it is just a rounding error - or minor changes in the values for g? (see this quote: we find that the force of earth’s gravity on your body at the equator is 9.798 m/s2 times the mass of your body, whereas at the poles it is 9.863 m/s2 times the mass of your body) There is evidently some room for manouevre

OpenStudy (mrnood):

THIS site ALSO suggest you need to account for the rotational forces, and ALSO calculates about 2N for that effect https://sciencequestionswithchris.wordpress.com/2014/01/07/do-i-weigh-less-on-the-equator-than-at-the-north-pole/

OpenStudy (anonymous):

it could be. I'll try the same thing with those values for g.

OpenStudy (anonymous):

With those new values for g I got 4.615 and it's still incorrect

OpenStudy (anonymous):

Are you saying I should subtract 2N?

OpenStudy (mrnood):

both effect act to reduce your weight at the equator so if the g effecct is 4N and the rotation effect is 2 N then the overall effect would be 6N less at the equator. (the site above quotes 1% - which would be 7N for the 71 kg body)

OpenStudy (mrnood):

I think you will struggle to get exactly the right answer - If it is online then it probably only going to accept your answer if it EXACTLY matches the programmed answer. You do not know what precision to enter it, or the values to take for g. I think you have shown you understand the concept so unless it is a compulsory test I should move on. Not all questions are 'good questions' for the sake of learning

OpenStudy (anonymous):

you need to calculate the gravity at the north pole first Fg= ( G m1 M2 )/ r^2 G :universal gravitation constant = 6.673 x 10^(-11) N m2/kg2 m1: mass of the earth m2: body mass ( man) but this mass also too small compare to the earth so it could be ignored r: earth radius

OpenStudy (anonymous):

Equatorial radius (km) 6378.1 Polar radius (km) 6356.8

OpenStudy (anonymous):

Core radius (km) 3485

OpenStudy (mrnood):

You cannot ignore m2 THAT is the mass of the body 71 The force is directly proportional to THAT However - you will still come up with 'debatable answers, and unless she is given the values ot use , and the precision to answer, then she is unlikely ot exactly match the 'computer' answer

OpenStudy (mrnood):

you have quoted radius to the nearest 100m - where was that measured - to the top of the hill? or the bottom of the valley?